Please help me with this problem. I am not quite sure how to set it up.
A physician invests $24,000 in two bonds. If one bond yields 6% and the other yields 12 %, how much is invested in each if the annual income from both bonds is $1980?

Answer by tutorcecilia(2152)

Investment x + Investment y = total invested
x+y=24,000
or x+(24,000-x)=24,000
.
ax+by=c
(Rate)(investment) + (Rate) (investment) = total interest earned
.06(x)+.12(24,000-x)=1980
.06x+2880-.12x=1980
.06x-.12x=-2880+1980 [solve for x]
-.06x=-900
-.06x/-.06=-900/-.06
x=15,000
.
So, if x = 15,000
Than y=(24,000-x)=(24,000-15,000)=9,000
Checking: x+y=24,000
15,000+9,000 = 24,000 [checks out]